# OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark for Vision Language Models

### v1

@@ 3.2.2 Question-Answer Annotation
We employ multiple-choice questions, including binary (true/false) and four-option formats, ensuring standardized evaluation and minimizing annotation errors. While adhering to structured formats, we prioritize natural and contextually rich language to prevent overly templated expressions.
Instead of rigid, mathematical phrasing (e.g., “Is [A] object on the right of [B] object?”), We use conversational styles more aligned with real-world reasoning (e.g., “If you’re entering the classroom, on which side are the students?”). This approach encourages models to rely on contextual and relative spatial references rather than predefined templates. To maintain clarity, we conduct multiple rounds of validation, resolving ambiguities in referents and ensuring unique answers. In particular, we ensure the quality of the data by means of cross-validation among different annotators.
We employ six human annotators for data labeling. All annotators receive comprehensive training to ensure a thorough understanding of the spatial reasoning tasks, enabling them to provide unbiased and discriminative assessments. All the annotators are from the authors of this work.
@@ 3.3 Improving Visual Spatial Reasoning Abilities
Given the diversity of spatial reasoning tasks in OmniSpatial, we further explore methods to enhance VLMs’ spatial reasoning capabilities.
3.3.1 PointGraph: Enhancing Spatial Reasoning via Point Relationships
We attempt to employ existing expert models, such as the Segment Anything Model [49], to partition images into distinct regions. This process extracts pixel clusters corresponding to multiple objects in the image, allowing models to estimate object centers and perform subsequent spatial reasoning tasks more effectively. As shown in Fig. 5, We attempt to enhance the spatial relationship comprehension of VLMs by explicitly incorporating additional spatial information cues into the models.
3.3.2 SpatialCoT: Enhancing Spatial Imagination via Novel View Synthesis
We observe that VLMs are largely constrained by textual reasoning paradigms, whereas humans inherently possess robust spatial imagination, commonly known as “mental imagery”. To some extent, this capability of spatial imagination corresponds to novel view synthesis of objects or scenes. Fortunately, recent advances in 3D generative models [100, 118, 87, 66] can perform such tasks effectively. In our work, we employ InstantMesh [118] to synthesize novel views from input images, providing additional visual cues to enhance spatial imagination.
Method Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
@@ 3.3.2 SpatialCoT: Enhancing Spatial Imagination via Novel View Synthesis
GPT-3.5-turbo [98] 30.67 - 38.38 29.19 38.35 28.76 36.91 0.82 24.00 42.16 33.67 35.90
GPT-4-turbo [77] 34.06 - 42.97 37.40 41.18 28.95 40.00 22.27 26.32 31.37 33.99 35.42
@@
GPT-4o-mini-2024-07-18 [45] 42.64 8 55.95 50.29 54.59 43.43 44.91 22.47 29.42 61.57 36.76 34.22
GPT-4o-2024-11-20 [45] 47.81 5 65.54 57.23 56.47 52.38 54.09 26.29 25.48 75.98 39.49 39.76
GPT-4.1-nano-2025-04-14 [78] 42.62 9 50.90 53.85 54.90 40.95 42.42 24.40 30.11 53.59 37.23 33.73
GPT-4.1-mini-2025-04-14 [78] 48.87 4 64.32 56.53 59.06 60.19 56.36 29.28 30.19 72.55 39.57 39.28
GPT-4.1-2025-04-14 [78] 51.78 2 66.22 64.74 60.00 65.33 60.18 31.75 30.06 70.98 40.64 39.04
Claude-3-5-sonnet-20241022 [3] 46.86 6 54.05 54.57 58.12 68.38 53.09 26.60 31.74 70.00 34.79 39.52
Claude-3-7-sonnet-20250219 [3] 47.53 5 57.57 55.95 56.71 63.81 59.09 29.48 28.39 72.16 36.06 36.63
Gemini-2.0-flash-lite-02-05 [2] 44.03 8 59.19 46.71 60.24 49.52 53.27 21.65 31.23 66.47 36.81 38.80
Gemini-2.0-flash-exp [2] 48.40 2 61.89 56.01 51.76 63.43 59.09 20.82 33.81 72.75 39.20 39.28
Gemini-2.5-flash-preview-05-20 [2] 52.12 1 67.57 62.72 68.24 73.33 60.91 38.14 34.19 75.49 35.90 33.73
@@
o1-2024-12-17 [46] 50.36 6 71.62 60.98 57.65 63.81 60.00 39.18 27.10 71.57 38.03 36.14
o4-mini-2025-04-16 [79] 52.77 3 72.97 59.83 60.00 73.33 61.82 34.02 36.77 73.53 40.69 40.96
o3-2025-04-16 [79] 56.33 1 71.89 66.18 61.18 68.57 65.45 40.21 29.68 77.06 48.40 48.19
Claude-3-7-sonnet-20250219-thinking [3] 48.62 7 57.21 59.73 53.73 67.94 57.27 30.24 28.17 68.63 37.94 36.95
Gemini-2.5-flash-05-20-thinking [2] 53.16 3 70.27 64.74 61.18 72.38 58.18 35.05 36.13 74.12 40.96 32.53
Gemini-2.5-pro-preview-05-06 [2] 55.19 2 67.57 71.39 62.35 75.24 64.55 43.30 34.84 74.51 38.03 37.35
@@
LLavA-1.5-vicuna-7B [64] 34.97 15 54.46 31.23 35.29 36.19 33.94 29.01 24.18 55.60 34.66 36.14
LLaVA-onevision-qwen2-7B [54] 35.68 14 43.24 38.15 32.94 29.52 41.82 28.87 22.58 47.06 36.17 37.35
LLaVA-onevision-qwen2-72B [54] 45.66 6 62.16 50.29 54.12 60.95 56.36 22.68 25.81 76.47 37.23 33.73
Gemma-3-4B [47] 39.79 11 41.89 49.71 56.47 27.62 36.36 23.71 24.52 59.80 36.17 38.55
Gemma-3-12B [47] 43.71 8 54.05 54.91 54.12 47.62 45.45 16.49 30.32 63.73 36.70 33.73
Gemma-3-27B [47] 44.75 7 56.76 55.78 57.65 50.48 52.73 27.84 29.03 64.71 33.51 32.53
InternVL3-2B [130] 37.98 13 50.00 40.58 43.29 40.00 40.55 21.86 28.52 55.49 35.11 33.01
InternVL3-8B [130] 41.60 9 52.43 40.87 48.94 51.05 44.77 24.95 28.63 64.20 38.62 40.96
InternVL3-14B [130] 45.94 5 54.32 60.17 50.35 51.81 51.45 28.04 28.26 68.04 35.37 34.46
InternVL3-38B [130] 48.48 2 63.42 63.58 54.59 58.29 50.55 29.90 28.52 72.16 36.76 33.49
InternVL3-78B [130] 49.33 1 63.78 63.12 56.24 59.24 51.45 27.63 30.19 74.51 38.46 35.90
Qwen-VL2.5-3B [115] 40.30 10 55.41 47.51 46.12 42.29 44.73 32.16 23.87 59.41 33.30 30.84
Qwen-VL2.5-7B [115] 39.18 12 58.38 35.09 50.12 45.33 44.00 31.13 29.42 64.51 33.19 37.35
Qwen-VL2.5-32B [115] 47.36 4 63.06 55.09 51.76 66.29 56.91 26.39 27.48 68.04 37.50 40.24
Qwen-VL2.5-72B [115] 47.85 3 58.38 60.12 50.12 59.81 53.64 26.19 33.03 71.37 36.81 36.39
@@
SpaceMantis-13B [15] 36.36 6 47.03 36.59 40.94 34.86 33.09 22.27 24.39 49.22 38.25 39.28
SpaceQwen2.5-VL-3B [15] 40.25 3 58.11 39.88 41.18 40.95 40.91 29.90 25.81 63.73 38.83 39.76
SpaceThinker-Qwen2.5VL-3B [15] 40.42 2 47.84 53.06 43.29 35.43 38.73 24.33 28.00 58.04 35.11 31.08
SpatialBot-3B [11] 35.68 6 43.24 38.15 32.94 29.52 41.82 28.87 22.58 47.06 36.17 37.35
RoboPoint-vicuna-v1.5-7B-lora [123] 35.85 6 57.03 28.61 34.82 37.33 40.55 29.90 22.71 50.20 38.72 40.96
RoboPoint-vicuna-v1.5-13B [123] 34.60 5 55.68 28.15 42.82 32.19 32.55 24.12 27.74 49.02 37.66 33.49
SoFar-Qwen2.5VL-3B [89] 45.14 1 56.49 51.16 54.12 53.14 52.73 31.75 22.88 71.60 36.56 41.69
@@
Human 92.63 - 96.53 97.30 92.94 97.14 94.55 91.30 87.63 99.02 95.74 93.98
Table 2: Evaluation on OmniSpatial. All models were tested 5 times and averaged to reduce randomness. Dark green indicates the best result within the group, and light green indicates the second-best result within the group. Gemini-2.5-Pro [97], InternVL3-78B [130], and SoFar [89] achieve the best performance within their respective groups. Among the various spatial intelligence categories, Complex Logic is the most challenging.
@@ 4.1 Evaluation Setup
To assess the spatial reasoning capabilities of existing VLMs, we conduct comprehensive experiments on OmniSpatial. This evaluation compares both proprietary and open-source models across various spatial tasks, using standardized benchmarks to ensure fairness and reproducibility.
Benchmark Models  We evaluate a range of state-of-the-art vision-language models, including:
  • Proprietary Models: These include GPT-4o series [45], GPT-4.1 series, Claude series [3], and Gemini series [2], which are state-of-the-art multimodal models accessed via an API under zero-shot settings with standardized prompts. All the system prompts can be found at  Appendix D.
  • Reasoning Models: We refer to current models employing internal long CoT as reasoning models. These models typically leverage reinforcement learning to enhance their reasoning capabilities, including models such as the o1, o3, o4-mini [46, 79], and Gemini-2.5 series [2, 97]. Due to the explicit nature of CoT, we utilize an additional text-based evaluation model [98] to determine the correctness of responses, similar to LLM-as-a-Judge [129].
  • Open Source Models: Our open-source set comprises Qwen-VL series [6], InternVL series [130], Gemma series [47] and LLaVA-Onevision series [63]. These models are currently the most advanced open-source VLMs. All models are deployed locally and use standardized prompts.
  • Specialized Spatial Models: These models, including SpatialVLM [15], RoboPoint [123], SpatialBot [11], and SoFar [89], are designed specifically for spatial reasoning tasks. SpatialVLM integrates metric spatial information through 3D data synthesis and scene segmentation. RoboPoint and SpatialBot emphasize the role of point affordance in spatial reasoning.
@@ 4.2 Evaluation Metrics
We evaluate the spatial reasoning capabilities of models by measuring their accuracy on multiple-choice questions. For standard open-source and closed-source models, we experiment with four approaches: enforcing direct output result, regular expression parsing, JSON format parsing and LLM-as-a-Judge [129]. For Reinforcement Learning-based reasoning models, due to the inherent challenges in enforcing structured output, we employ GPT-4.1-mini to assess the correctness of the model’s responses by comparing them with the ground truth. More detailed ablation studies are discussed in the Section C.1.
Method Avg. Improve Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
@@
Table 3: Camparison of textual Chain-of-Thought and PointGraph on OmniSpatial.
@@ 4.3 Main Results
Overall Model Performance  As illustrated in Table 2, we observe the following findings: (i) Proprietary reasoning models, such as ChatGPT o3 [79] and Gemini-2.5-pro [2], achieve the highest performance, surpassing a  56% overall success rate; however, there remains a significant gap compared to human-level understanding, and require a lot of inference time and tokens. (ii) Open-source models also demonstrate competitive results, with large-scale models like InternVL3-78B [130] and Qwen-VL2.5-72B [115] achieving comparable performance to GPT-4.1-mini and Gemini-2.0-flash-exp. (iii) Specialized Spatial Reasoning Models, due to limitations in dataset coverage and model capacity, struggle to achieve substantial improvements on comprehensive benchmarks.
Category-wise Analysis  Breaking down the results, we observe notable performance differences across spatial reasoning categories: (i) Leveraging their extensive world knowledge and local understanding capabilities, proprietary models have demonstrated strong performance in Dynamic Reasoning and Spatial Interaction, indicating that reasoning models possess high proficiency in temporal understanding, spatial relationship analysis, and map-based comprehension. (ii) For Pattern Geometric and Recognition Reasoning, which involve spatial imagination in planar geometry, even reasoning models designed for extended thinking can only achieve an accuracy of around 30% to 40%, slightly surpassing the random baseline. (iii) Current models exhibit limited Perspective Taking abilities, predominantly analyzing scenarios from an ego-centric viewpoint while struggling to imagine perspectives from others’ viewpoints.
@@
Qwen-VL2.5-3B
(w/ Zero-shot CoT) 40.89 - 59.61 36.54 37.59
@@
Impact of PointGraph & Spatial CoT  To test whether structured segmentation improves performance, we apply PointGraph as a preprocessing step for GPT-4.1, Gemini-2.5-flash and Qwen-VL2.5-7B. Results in Table 3 show a clear accuracy boost, particularly in the Dynamic Reasoning and Perspective-Taking Track, validating the benefits of integrating structured object representation, while traditional textual CoT difficult to bring about significant improvement. Fig. 5 further demonstrates the effectiveness of our proposed Spatial CoT on the OmniSpatial Perspective-Taking track. Through novel view synthesis facilitated by InstantMesh, both GPT-4.1 and Qwen-VL2.5-7B exhibit significant performance improvements, validating the effectiveness of explicit spatial imagination.

### v2

@@ 3.2.2 Question-Answer Annotation
We design multiple-choice questions, including binary (true/false) and four-option formats, to enable standardized evaluation while minimizing annotation bias. To ensure naturalness, questions are phrased in conversational and context-rich styles (e.g., “If you are entering the classroom, on which side are the students?”) rather than rigid templates (e.g., “Is [A] on the right of [B]?”). This encourages models to rely on contextual and relational reasoning rather than memorized patterns.
To guarantee clarity and answer uniqueness, we perform multiple rounds of validation and resolve ambiguous references through cross-checking among annotators. Six trained annotators—initially the authors—were involved in labeling. Each annotator underwent task-specific training, and annotations were cross-validated to reduce subjectivity. The inter-annotator agreement reached Krippendorff’s $\alpha=0.84$, indicating a high level of consistency.
@@ 3.3 Improving Visual Spatial Reasoning Abilities
3.3.1 PointGraph: Explicit Modeling of Object Relationships
To strengthen models’ understanding of spatial relations, we introduce PointGraph, which constructs structured representations of object relationships within an image. Concretely, we employ open-vocabulary grounding models such as Florence-2 to localize multiple objects and extract their centers and bounding boxes. These detections are then assembled into a JSON-style scene graph encoding object identities and relative positions. By concatenating this structured spatial description with the original query, VLMs are provided with explicit geometric cues that facilitate more accurate reasoning about distances, directions, and configurations (see Fig.˜5).
3.3.2 SpatialCoT: Stimulating Spatial Imagination via Novel Views
Human spatial reasoning often relies on mental imagery—the ability to imagine how a scene looks from different viewpoints. Inspired by this, we design SpatialCoT, which augments visual inputs with 3D novel views to enrich spatial imagination (Shi et al., 2023, Qi et al., 2023b, Liu et al., 2024d). Specifically, we adopt InstantMesh (Xu et al., 2024a) to synthesize six additional perspectives for each input image and compose them into a multi-view collage. This collage is then fed, along with the question, into the VLM as part of a chain-of-thought prompting pipeline. The additional perspectives provide strong geometric priors, helping models disambiguate occlusions, perspective-taking, and other view-dependent reasoning tasks.
Table 2: Evaluation on OmniSpatial-test. All models were tested 5 times and averaged to reduce randomness. Dark green indicates the best result and light green indicates the second-best result within the group. Gemini-2.5-Pro (Reid et al., 2024), InternVL3-78B (Zhu et al., 2025), and SoFar (Qi et al., 2025) achieve the best performance within their respective groups.
Method
Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
@@ 3.3.2 SpatialCoT: Stimulating Spatial Imagination via Novel Views
GPT-3.5-turbo (Roumeliotis & Tselikas, 2023) 30.67 - 38.38 29.19 38.35 28.76 36.91 0.82 24.00 42.16 33.67 35.90
GPT-4-turbo (OpenAI, 2023) 34.06 - 42.97 37.40 41.18 28.95 40.00 22.27 26.32 31.37 33.99 35.42
@@
GPT-4o-mini-2024-07-18 (Hurst et al., 2024) 42.64 8 55.95 50.29 54.59 43.43 44.91 22.47 29.42 61.57 36.76 34.22
GPT-4o-2024-11-20 (Hurst et al., 2024) 47.81 5 65.54 57.23 56.47 52.38 54.09 26.29 25.48 75.98 39.49 39.76
GPT-4.1-nano-2025-04-14 (OpenAI, 2025a) 42.62 9 50.90 53.85 54.90 40.95 42.42 24.40 30.11 53.59 37.23 33.73
GPT-4.1-mini-2025-04-14 (OpenAI, 2025a) 48.87 4 64.32 56.53 59.06 60.19 56.36 29.28 30.19 72.55 39.57 39.28
GPT-4.1-2025-04-14 (OpenAI, 2025a) 51.78 2 66.22 64.74 60.00 65.33 60.18 31.75 30.06 70.98 40.64 39.04
Claude-3-5-sonnet-20241022 (Anthropic, 2024) 46.86 6 54.05 54.57 58.12 68.38 53.09 26.60 31.74 70.00 34.79 39.52
Claude-3-7-sonnet-20250219 (Anthropic, 2024) 47.53 5 57.57 55.95 56.71 63.81 59.09 29.48 28.39 72.16 36.06 36.63
Gemini-2.0-flash-lite-02-05 (Anil et al., 2023) 44.03 8 59.19 46.71 60.24 49.52 53.27 21.65 31.23 66.47 36.81 38.80
Gemini-2.0-flash-exp (Anil et al., 2023) 48.40 2 61.89 56.01 51.76 63.43 59.09 20.82 33.81 72.75 39.20 39.28
Gemini-2.5-flash-preview-05-20 (Anil et al., 2023) 52.12 1 67.57 62.72 68.24 73.33 60.91 38.14 34.19 75.49 35.90 33.73
@@
o1-2024-12-17 (Jaech et al., 2024) 50.36 6 71.62 60.98 57.65 63.81 60.00 39.18 27.10 71.57 38.03 36.14
o4-mini-2025-04-16 (OpenAI, 2025b) 52.77 3 72.97 59.83 60.00 73.33 61.82 34.02 36.77 73.53 40.69 40.96
o3-2025-04-16 (OpenAI, 2025b) 56.33 1 71.89 66.18 61.18 68.57 65.45 40.21 29.68 77.06 48.40 48.19
Claude-3-7-sonnet-20250219-thinking (Anthropic, 2024) 48.62 7 57.21 59.73 53.73 67.94 57.27 30.24 28.17 68.63 37.94 36.95
Gemini-2.5-flash-05-20-thinking (Anil et al., 2023) 53.16 3 70.27 64.74 61.18 72.38 58.18 35.05 36.13 74.12 40.96 32.53
Gemini-2.5-pro-preview-05-06 (Anil et al., 2023) 55.19 2 67.57 71.39 62.35 75.24 64.55 43.30 34.84 74.51 38.03 37.35
@@
LLavA-1.5-vicuna-7B (Liu et al., 2024c) 34.97 15 54.46 31.23 35.29 36.19 33.94 29.01 24.18 55.60 34.66 36.14
LLaVA-onevision-qwen2-7B (Li et al., 2024a) 35.68 14 43.24 38.15 32.94 29.52 41.82 28.87 22.58 47.06 36.17 37.35
LLaVA-onevision-qwen2-72B (Li et al., 2024a) 45.66 6 62.16 50.29 54.12 60.95 56.36 22.68 25.81 76.47 37.23 33.73
Gemma-3-4B (Kamath et al., 2025) 39.79 11 41.89 49.71 56.47 27.62 36.36 23.71 24.52 59.80 36.17 38.55
Gemma-3-12B (Kamath et al., 2025) 43.71 8 54.05 54.91 54.12 47.62 45.45 16.49 30.32 63.73 36.70 33.73
Gemma-3-27B (Kamath et al., 2025) 44.75 7 56.76 55.78 57.65 50.48 52.73 27.84 29.03 64.71 33.51 32.53
InternVL3-2B (Zhu et al., 2025) 37.98 13 50.00 40.58 43.29 40.00 40.55 21.86 28.52 55.49 35.11 33.01
InternVL3-8B (Zhu et al., 2025) 41.60 9 52.43 40.87 48.94 51.05 44.77 24.95 28.63 64.20 38.62 40.96
InternVL3-14B (Zhu et al., 2025) 45.94 5 54.32 60.17 50.35 51.81 51.45 28.04 28.26 68.04 35.37 34.46
InternVL3-38B (Zhu et al., 2025) 48.48 2 63.42 63.58 54.59 58.29 50.55 29.90 28.52 72.16 36.76 33.49
InternVL3-78B (Zhu et al., 2025) 49.33 1 63.78 63.12 56.24 59.24 51.45 27.63 30.19 74.51 38.46 35.90
Qwen-VL2.5-3B (Wang et al., 2024b) 40.30 10 55.41 47.51 46.12 42.29 44.73 32.16 23.87 59.41 33.30 30.84
Qwen-VL2.5-7B (Wang et al., 2024b) 39.18 12 58.38 35.09 50.12 45.33 44.00 31.13 29.42 64.51 33.19 37.35
Qwen-VL2.5-32B (Wang et al., 2024b) 47.36 4 63.06 55.09 51.76 66.29 56.91 26.39 27.48 68.04 37.50 40.24
Qwen-VL2.5-72B (Wang et al., 2024b) 47.85 3 58.38 60.12 50.12 59.81 53.64 26.19 33.03 71.37 36.81 36.39
@@
SpaceMantis-13B (Chen et al., 2024) 36.36 6 47.03 36.59 40.94 34.86 33.09 22.27 24.39 49.22 38.25 39.28
SpaceQwen2.5-VL-3B (Chen et al., 2024) 40.25 3 58.11 39.88 41.18 40.95 40.91 29.90 25.81 63.73 38.83 39.76
SpaceThinker-Qwen2.5VL-3B (Chen et al., 2024) 40.42 2 47.84 53.06 43.29 35.43 38.73 24.33 28.00 58.04 35.11 31.08
SpatialBot-3B (Cai et al., 2024) 35.68 6 43.24 38.15 32.94 29.52 41.82 28.87 22.58 47.06 36.17 37.35
RoboPoint-vicuna-v1.5-7B-lora (Yuan et al., 2024) 35.85 6 57.03 28.61 34.82 37.33 40.55 29.90 22.71 50.20 38.72 40.96
RoboPoint-vicuna-v1.5-13B (Yuan et al., 2024) 34.60 5 55.68 28.15 42.82 32.19 32.55 24.12 27.74 49.02 37.66 33.49
SoFar-Qwen2.5VL-3B (Qi et al., 2025) 45.14 1 56.49 51.16 54.12 53.14 52.73 31.75 22.88 71.60 36.56 41.69
@@
Human 92.63 - 94.62 96.07 91.38 95.11 92.15 89.02 85.90 98.53 94.30 90.26
@@ 4.1 Evaluation Setup
To systematically assess spatial reasoning, we evaluate models on the OmniSpatial benchmark under a unified protocol. Our evaluation covers both proprietary and open-source systems, spanning general-purpose VLMs, reasoning-oriented LLMs, and spatially specialized models, thereby ensuring broad coverage and fair comparison. We consider four groups of state-of-the-art models:
  • Proprietary Models. These include the GPT-4o and GPT-4.1 families (Hurst et al., 2024), the Claude series (Anthropic, 2024), and Gemini series (Anil et al., 2023). All are accessed through APIs under zero-shot settings with standardized system prompts (see Appendix˜D).
  • Reasoning Models. We categorize models that explicitly employ long chain-of-thought reasoning, often enhanced through reinforcement learning, such as o1, o3, o4-mini (Jaech et al., 2024, OpenAI, 2025b), and Gemini-2.5 (Anil et al., 2023, Reid et al., 2024). Because their outputs are less amenable to strict parsing, we additionally rely on an automatic judge to compare answers against ground truth, following the LLM-as-a-Judge paradigm (Zheng et al., 2023).
  • Open-Source Models. This set includes Qwen-VL (Bai et al., 2023), InternVL (Zhu et al., 2025), Gemma (Kamath et al., 2025), and LLaVA-OneVision (Liu et al., 2023b). These are locally deployed with standardized prompts to ensure reproducibility.
  • Specialized Spatial Models. We further benchmark models designed specifically for spatial reasoning, including SpatialVLM (Chen et al., 2024), RoboPoint (Yuan et al., 2024), SpatialBot (Cai et al., 2024), and SoFar (Qi et al., 2025). These systems incorporate explicit spatial signals such as metric 3D information, point affordances, or semantic orientation to improve reasoning.
@@ 4.2 Evaluation Metrics
We measure accuracy on multiple-choice questions. For standard proprietary and open-source models, we test four output protocols: direct answer, regular-expression parsing, JSON parsing, and LLM-as-a-Judge (Zheng et al., 2023). For reasoning-oriented models with unstructured CoT outputs, correctness is assessed by GPT-4.1-mini against ground truth. Ablations of these evaluation strategies are provided in Section˜C.1.
Table 3: Camparison of textual Chain-of-Thought and PointGraph on OmniSpatial-test.
Method
Avg. Improve Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
@@ 4.3 Main Results
Overall Model Performance  As illustrated in Table˜2, we observe the following findings: (i) Proprietary reasoning models, such as ChatGPT o3 (OpenAI, 2025b) and Gemini-2.5-pro (Anil et al., 2023), achieve the highest performance, surpassing a  56% overall success rate; however, there remains a significant gap compared to human-level understanding, and they require a lot of inference time and tokens. (ii) Open-source models also demonstrate competitive results, with large-scale models like InternVL3-78B (Zhu et al., 2025) and Qwen-VL2.5-72B (Wang et al., 2024b) achieving comparable performance to GPT-4.1-mini and Gemini-2.0-flash-exp. (iii) Specialized Spatial Reasoning Models, due to limitations in dataset coverage and model capacity, struggle to achieve substantial improvements on comprehensive benchmarks.
Category-wise Analysis  We observe notable performance differences across spatial reasoning categories: (i) Leveraging their extensive world knowledge and local understanding capabilities, proprietary models have demonstrated strong performance in Dynamic Reasoning and Spatial Interaction, indicating that reasoning models possess high proficiency in temporal understanding, spatial relationship analysis, and map-based comprehension. (ii) For Pattern Geometric Reasoning, which involves spatial imagination in planar geometry, even reasoning models designed for extended thinking can only achieve an accuracy of around 30% to 40%, slightly surpassing the random baseline. (iii) Current models exhibit limited perspective-taking abilities, predominantly analyzing scenarios from an ego-centric viewpoint while struggling to imagine perspectives from others’ viewpoints.
@@
Qwen‑VL2.5‑3B
(w/ Zero‑shot CoT) 40.89 - 59.61 36.54 37.59
@@
Impact of PointGraph & Spatial CoT  To test whether structured segmentation improves performance, we apply PointGraph as a preprocessing step for GPT-4.1, Gemini-2.5-flash and Qwen-VL2.5-7B. Results in Table˜3 show a clear accuracy boost, particularly in the Dynamic Reasoning and Perspective-Taking Track, validating the benefits of integrating structured object representation, while traditional textual CoT difficult to bring about significant improvement. Fig.˜5 further demonstrates the effectiveness of our proposed Spatial CoT on the OmniSpatial Perspective-Taking track. Through novel view synthesis facilitated by InstantMesh, both GPT-4.1 and Qwen-VL2.5-7B exhibit significant performance improvements, validating the effectiveness of explicit spatial imagination.
4.4 Training Exploration
Supervised Training on OmniSpatial-train.  We further study whether OmniSpatial-train can effectively teach spatial skills instead of merely fitting templates. Starting from a strong open-source baseline, supervised fine-tuning on 6.9K samples yields a substantial +7.82 point average gain over zero-shot, with consistent improvements across dynamic, interaction, and perspective-taking oriented tracks. In contrast, training on a much larger 200K template-style corpus that follows the construction process of VSI-Bench (Yang et al., 2024) brings only a marginal +1.29 average gain, underscoring the value of diverse, manually curated spatial tasks over synthetic templates.
Table 5: Training Exploration on OmniSpatial-test.
Method Avg. Improve
Manipulate
Motion
Analysis
Traffic
Analysis
Locate
Geospatial
Strategy
Pattern
Recognition
Geometric
Reasoning
Ego
Centric
Allo
Centric
Hypothetical
Qwen-VL2.5-3B - - - - - - - - - - - -
Zero-shot 40.30 - 55.41 47.51 46.12 42.29 44.73 32.16 23.87 59.41 33.30 30.84
+ OmniSpatial-train (6.9K) 48.12 +7.82 59.19 51.16 50.12 51.81 45.45 34.02 31.74 68.63 36.76 35.90
+ Template corpus (200K) 41.59 +1.29 50.90 46.71 40.94 45.33 40.55 28.87 24.39 72.55 35.11 33.01
Generalization to VSI-Bench.  To examine cross-benchmark generalization, we adopt the SpaceR-7B pipeline and compare supervised training with and without OmniSpatial. Adding OmniSpatial improves the overall score on VSI-Bench from 41.68 to 43.68. Notably, it boosts categories requiring ordering, counting, and metric/room-size reasoning (e.g., appearance_order, obj_counting, obj/room_size). These results indicate that OmniSpatial provides complementary supervision that transfers to external spatial tasks rather than overfitting to in-benchmark patterns.
Table 6: Generalization on VSI-Bench. Training with OmniSpatial yields consistent gains.
Method overall
Appearance
Order
Obj Abs
Distance
Obj
Counting
Obj Rel
Distance
Obj
Size
Room
Size
Route
Planning
Obj Rel
Direction
Qwen-VL2.5-3B - - - - - - - - -
Zero-shot 34.06 34.63 11.39 40.18 36.20 46.11 38.58 31.96 36.92
+ SpaceR-7B 41.68 46.60 19.63 55.27 34.23 58.32 35.17 34.54 42.84
+ OmniSpatial 43.68 58.25 15.13 57.36 34.37 60.99 41.11 34.02 44.16